If z1, z2, z3 are complex numbers such that |z1|=|z2|=|z3|=   | (1/z1) + (1/z2) + (1/z3) | =1 then |z1+z2+z3|=?? How to solve these types of sums??

Grade:

2 Answers

Arun
5 years ago
Given : | z₁ | = | z₂ | = | z₃ | = 1 ............. (1) 

. . . . . . | 1/z₁ + 1/z₂ + 1/z₃ | = 1 ............ (2) 
__________________________ 

To Find : | z₁ + z₂ + z₃ |. 
__________________________ 

We note that 

(i) | z | = | z' |, ... z' is conjugate of z 

(ii) z·z' = | z |² = | z' |² 

(iii) Σ(z') = [ Σ(z) ]'. 
__________________________ 

If | z | = 1, then : z·z' = 1, i.e., (1/z) = z'...... (3) 

∴ | z₁ + z₂ + z₃ | 

= | ( z₁ + z₂ + z₃ )' | 

= | (z₁)' + (z₂)' + (z₃)' | 

= | 1/z₁ + 1/z₂ + 1/z₃ | ........ from (3) 

= 1 ...................................... from (2) 
__________________ 

∴ | z₁ + z₂ + z₃ | = 1 
ankit singh
3 years ago
Given : | z₁ | = | z₂ | = | z₃ | = 1 ............. (1) 
. . . . . . | 1/z₁ + 1/z₂ + 1/z₃ | = 1 ............ (2) 
__________________________ 
To Find : | z₁ + z₂ + z₃ |. 
__________________________ 
We note that 
(i) | z | = | z' |, ... z' is conjugate of z 
(ii) z·z' = | z |² = | z' |² 
(iii) Σ(z') = [ Σ(z) ]'. 
__________________________ 
If | z | = 1, then : z·z' = 1, i.e., (1/z) = z'...... (3) 
∴ | z₁ + z₂ + z₃ | 
= | ( z₁ + z₂ + z₃ )' | 
= | (z₁)' + (z₂)' + (z₃)' | 
= | 1/z₁ + 1/z₂ + 1/z₃ | ........ from (3) 
= 1 ...................................... from (2) 
__________________ 
∴ | z₁ + z₂ + z₃ | = 1